OpenStudy (anonymous):

Find centripetal acceleration For a body in a motion with, v = 2t i + t^2 j i --- x direction j ---- y direction.

OpenStudy (anonymous):

This is trajectory actually a parabola.I don't think that the particle executes circular motion.

OpenStudy (anonymous):

Nevertheless, for any curvilinear motion, a centripetal acceleration always exists.

OpenStudy (anonymous):

Well in that case at what time do you want the centripetal acceleration to be calculated or is it a general case that you want?

OpenStudy (anonymous):

A general way to attempt this problem.

OpenStudy (experimentx):

do you know the answer?

OpenStudy (anonymous):

Not yet.

OpenStudy (anonymous):

What is the origin with respect to which you want to calculate the centripetal acceleration.

OpenStudy (anonymous):

I mean the center.

OpenStudy (anonymous):

can u find magnitude of v

OpenStudy (anonymous):

Yes it is the magnitude of the velocity vector.

OpenStudy (experimentx):

I'm not sure .... most likely, you need radius, x = t^2, y=t^3/3 and you need transverse component of velocity ... (most likely, this is related to curvature problem)

OpenStudy (anonymous):

@kartiksriramk I don't follow. The equation of the velocity should suffice.

OpenStudy (anonymous):

I dont think this will work but,